Discuss with Askiitians Tutors

A man lends $ 1000 in four sums. If he gets 3% for $ 200, 3 1/2% for $400 and 5% for $ 250, ;what percent must he get for the remainder, if his average interest is 3.85% ?


Indira pedaprolu
13 Years agoGrade

1 Answer

aa aa
13 Years ago

3% for 200 => $6

3.5% for 400 => $14

5% for 250 => $12.50

Total interest = 3.85% of 1000 => 38.5

Remaining interest = 38.5 - 6 - 14 -12.5 = $6

Rate = 6/150*100 = 4%

Ans 4%